OpenStudy (anonymous):

put in radical form x^2/3

OpenStudy (anonymous):

\[x ^{3/2}\]

OpenStudy (anonymous):

dont look at the one in the question i asked, i messed it up

OpenStudy (anonymous):

\[\sqrt[3]{x^2}\]

OpenStudy (anonymous):

\[\sqrt[3]{x ^{2}}\]

OpenStudy (anonymous):

\[x^{\frac{p}{q}}=\sqrt[q]{x^p}\]

OpenStudy (anonymous):

think he wants sqrt of x^3...

OpenStudy (anonymous):

\[\sqrt{x^3}\]

OpenStudy (anonymous):

He switched the 2 and 3 after he posted...